OpenStudy (anonymous):

A system of linear inequalities is shown below. x - y > 3 y + x ≤ 2 Describe the steps to graph the solution set to the system of inequalities.

OpenStudy (anonymous):

The first eqaution gives you y > x - 3....So graph the line y = x - 3 and shade the region where y is greater than that line. The second eqaution gives you y < or = to -x + 2...So draw a dotted (because or equal to) line y = -x +2 and shade the region where y is less than. Ok now the region that is shade from both eqautions satisfy both ineqaulities and you can distinctivly shade that as your final answer

OpenStudy (campbell_st):

well I always thought > or < needed a dotted line and >= or <= needed a solid line...

OpenStudy (campbell_st):

I'd say, for the intercepts for both lines plot the intercepts draw the appropriate line, solid or dotted shade the region.... testing with a point like (1,1) to check the common region will contain both shadings